A 90% confidence interval around the variance was calculated. For n=27, the confidence interval was 35.0-88.4. Give the variance of the sample.
Answer by stanbon(75887)

A 90% confidence interval around the variance was calculated. For n=27, the confidence interval was 35.0-88.4. Give the variance of the sample.
-----------------------------
(n-1)s^2/(Ch-Sq right) = 35
26*s^2/41.923 = 35
s^2 = 35*41.923/26
s^2 = 56.434
--------------------
(n-1)s^2/(Ch-Sq left) = 88.4
26*s^2/15.379 = 88.4
s^2 = 88.4*15.379/26
s^2 = 52.2886
------------------------
The discrepancy is probably due to my Ch-Sq table truncating its values.
It gives the Ch-Sq. left and right values as indicated. Your chart
may give values with greater decimal precision.
-------------------------
Cheers,
Stan H.